1. Current travels clockwise through a loop. Sketch the direction of the magnetic field around both sides of the loop (use \odot(out of page), \otimes(into the page)).

2. Two wires carry 3 A in opposite directions, as shown. Calculate the magnitude and direction of the magnetic field at each point.

(1) Point a:

(2)Point b (directly between wires)

  (3) Point c:
Answer #1

(a)

(2)

at point a

B = uoi/ 2 pi a - uo i / 2 pi ( a+ b)

= uo i/ 2 pi ( 1/a- 1/(a+ b)

= 2 * 10^-7 ( 3 A) ( 1/5 - 1/8)

=0.45 * 10^-7 T

(2)

at point b

B = uoi/ 2 pi (b/2) - uo i / 2 pi (b/2)

= 0 T

(3)

at point c

B = uoi/ 2 pi c - uo i / 2 pi ( b+c)

= uo i/ 2 pi ( 1/c- 1/(b+c)

= 2 * 10^-7 ( 3) ( 1/2- 1/5)

=1.8 * 10^-7 T